OpenStudy (anonymous):

|x+1|-4=-1 Can someone explain this to me? I got 2 and 4 for my answers but the correct answers are 2 and -4. I don't get it.

jhonyy9 (jhonyy9):

I x+1 I =3

jhonyy9 (jhonyy9):

so x+1=3 so x=2 and x+1=-3 so x= -4

OpenStudy (anonymous):

do i have to rewrite it that way? i just removed the bars and worked the problem and got the 2

OpenStudy (anonymous):

x+1 = 3 x = 2 or x + 1 = -3 x = -4

OpenStudy (anonymous):

I have to isolate what's between the bars on one side?

OpenStudy (anonymous):

How come it's equal to 3??? Adding 4 to both sides makes it equal to 5. I don't understand.

OpenStudy (phi):

are you asking why 4+ -1 = 3?

OpenStudy (anonymous):

\[|x+1|-4=-1\] add 4 \[|x+1|=3\] the solve two equations separately \[x+1=3\] or \[x+1=-3\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

@ phi, i had it written as equal to 1, my mistake, sorry